# SplatTalk: 3D VQA with Gaussian Splatting

### v1

@@ Abstract
Language-guided 3D scene understanding is important for advancing applications in robotics, AR/VR, and human-computer interaction, enabling models to comprehend and interact with 3D environments through natural language. While 2D vision-language models (VLMs) have achieved remarkable success in 2D VQA tasks, progress in the 3D domain has been significantly slower due to the complexity of 3D data and the high cost of manual annotations. In this work, we introduce SplatTalk, a novel method that uses a generalizable 3D Gaussian Splatting (3DGS) framework to produce 3D tokens suitable for direct input into a pretrained LLM, enabling effective zero-shot 3D visual question answering (3D VQA) for scenes with only posed images. During experiments on multiple benchmarks, our approach outperforms both 3D models trained specifically for the task and previous 2D-LMM-based models utilizing only images (our setting), while achieving competitive performance with state-of-the-art 3D LMMs that additionally utilize 3D inputs.
@@ 1 Introduction
Previous works have explored integrating semantic features into NeRF [21] or 3D Gaussian Splatting (3DGS) [38], which rely only upon RGB image inputs. However, research on effectively associating 3D geometry with language fetures remains limited. Some methods align individual 3D locations with language features via implicit functions optimized individually for every scene [21, 10]. Others employ SAM [22] to first segment objects and object parts, followed by CLIP [39] to extract object-centric features from the generated masks. These approaches allows models to query information about individual points and/or objects. However, it disregards spatial relationships between objects, making it unsuitable for 3D VQA spatial reasoning tasks.
@@ 2 Related Work
Generalizable 3D Gaussian Splatting. This body of work explores generalizable 3DGS as an alternative to the per-scene optimization paradigm used in traditional 3DGS methods, aiming to improve novel view synthesis across unseen scenes. FreeSplat [46] introduces a cost-volume-based approach to predict 3D Gaussian parameters from sparse-view inputs, improving reconstruction quality without requiring per-scene optimization. One of the key contributions of this work is the introduction of 3D Gaussian triplets, which are latent feature vectors that serve as a basis representation of the scene, capturing the scene information in a compact form. The method then employs a Pixel-wise Triplet Fusion module, which leverages a lightweight GRU network to fuse overlapping 3D Gaussian triplets from multiple views. This process significantly reduces 3D Gaussian redundancy.
Similarly, [8, 5] leverage multi-view stereo (MVS) cues to refine depth estimation and improve geometric consistency in 3D Gaussian representations. These methods primarily focus on RGB novel view synthesis from sparse inputs (typically views). Our work builds on the SOTA model, FreeSplat [46], but diverges significantly in both objectives and methodology. First, while previous methods focus on photorealistic RGB reconstruction, we aim at semantic 3D scene understanding. Second, instead of reconstructing only a localized portion of the scene as shown in these works, our goal is to represent entire 3D environments, capturing both spatial and semantic relationships holistically.
@@
ChatSplat [6] is a concurrent work that explores natural language conversation within the 3DGS framework. However, it optimizes on a per-scene basis and requires rendering 3D features into 2D feature maps during inference. Additionally, its training pipeline is not fully end-to-end, as it first trains the RGB Gaussians separately before incorporating language features. In contrast, our approach leverages scene latent features that encapsulate both RGB and semantic information, enabling the joint optimization of both modalities in a fully end-to-end manner.
@@ 3 Method
We provide an overview of our approach in Fig. 2, which consists of three main steps: 1) Training the feature autoencoder, 2) Self-supervised 3D-Language Gaussian Splatting Training and 3) 3D VQA Inference. First, we extract the high-dimensional language features from LLaVA-OV [23] for each RGB image by using the pretrained image encoder and multimodal projector. Since these features are high-dimensional, unbounded and sparse, we train an autoconder to project them into a more compact, low-dimensional space. These low-dimensional features then serve as pseudo ground-truth for training the language representations in SplatTalk. SplatTalk architecture is a feed-forward network consists of a Gaussian encoder and decoder that predicts Gaussian parameters for rendering. During 3D VQA inference stage, we extract the trained language features from the 3D Gaussians and directly feed them into the LLM, along with the language prompt, to generate free-form responses.
@@ 3.1 Integrating Language Features
| SplatTalk | I | 61.7 | 14.2 | 32.7 | 17.1 | 32.2 | 26.1 | 41.3 |
| SplatTalk-ScanQA-FT | I | 77.1 | 15.4 | 38.7 | 22.3 | 38.3 | 42.0 | 44.7 |
| SplatTalk-3DVQA-FT | I | 77.5 | 15.6 | 38.5 | 22.4 | 38.3 | 47.6 | 49.4 |
3.2 Extracting 3D Language Features
During inference, we directly extract language features at the mean position of each 3D Gaussian. Although the rendering process weights features according to the Gaussians’ covariances and opacities, we demonstrate that the semantic feature derived from each Gaussian’s mean effectively captures the scene’s semantics. We further validate this choice by highlighting the conceptual similarity between our training optimization process and the Expectation-Maximization (EM) algorithm. For simplification, let be the loss, where is the predicted rendered feature map of frame at pixel and is the pseudo ground-truth feature map obtained from LLaVA-OV. Replacing by the rendering equation and solving for the optimal :
where . Note that this iterative optimization procedure, rendering and minimizing reconstruction errors, is conceptually analogous to the EM algorithm. Specifically, the procedure consists of: (1) an E-step, which estimates each Gaussian’s contribution to pixels across viewpoints given current Gaussian parameters, and (2) an M-step, which updates the Gaussian parameters to minimize reconstruction errors. Consequently, the optimal 3D feature for each Gaussian can be seen as a weighted sum of the 2D feature maps, with the weights representing each Gaussian’s contributions to the pixels, as computed through the rendering process. Therefore the set of 3D Gaussian features is expected to effectively encode the semantic content of the scene as whole from 2D inputs. The extracted 3D language features are then projected back into the original language embedding space using the pretrained decoder and directly provided as inputs to the LLM for downstream 3D-language tasks.
Note that in open-vocabulary segmentation task, querying features at Gaussian centers appear straightforward [4, 38, 48] because the task inherently assumes that each Gaussian corresponds to a distinct, semantically meaningful object or region. These approaches leverage pre-trained SAM-masked CLIP or LSeg embeddings, where features are object-centric, making it possible to retrieve objects simply by computing cosine similarity between a Gaussian’s feature and a text query. However, this assumption does not hold in 3D reasoning tasks. In our case, a single Gaussian does not necessarily encode an explicit object category. Instead, Gaussians contribute to a scene-level representation, where meaning can emerge from the spatial and relational structure of multiple features rather than a direct one-to-one mapping with objects. This makes querying individual Gaussians insufficient for tasks that require context-dependent inference rather than segmentation-based retrieval. Therefore, while our method queries at the mean Gaussians, it is not inherently intuitive or trivial to do so. Our proof using a theoretical connection to the EM algorithm in this section has provided a principled justification for this approach.
@@ 3.2 Extracting 3D Language Features
| SplatTalk | T | I | zero-shot | 19.6 | 60.3 | 44.0 | 35.8 | 35.5 | 61.8 | 41.8 |
| SplatTalk-ScanQA-FT | T | I | zero-shot | 28.9 | 66.0 | 43.2 | 28.3 | 33.8 | 60.0 | 41.5 |
| SplatTalk-3DVQA-FT | T | I | zero-shot | 24.3 | 57.0 | 36.1 | 7.8 | 16.4 | 35.0 | 26.5 |
3.3 Entropy Adaptive Gaussian Sampling
SplatTalk leverages FreeSplat [46]’s Gaussian triplet fusion module that significantly reduces the number of 3D Gaussians needed to represent a scene. However, effectively capturing fine details in complex scenes still requires a large number of Gaussians, often exceeding the context window size available for visual tokens in the LLM. Further, not all Gaussians are equally important for downstream 3D-language tasks. To effectively capture the most semantically informative aspects of the scene, we introduce an entropy-based adaptive Gaussian sampling strategy. Specifically, we select the top Gaussians whose language features exhibit the highest entropy, where matches the number of visual tokens accepted by the LLM.
@@ 4.1 Implementation Details
| Sampling Method | EM@1 | EM@1-R | METEOR | ROUGE | CIDEr | SPICE |
| Random | 18.6 | 33.5 | 13.5 | 31.7 | 59.9 | 13.0 |
| Point Density | 17.9 | 34.4 | 14.2 | 32.1 | 60.8 | 14.8 |
| FPS | 18.9 | 34.8 | 14.4 | 33.0 | 62.7 | 14.2 |
| Entropy | 19.8 | 34.0 | 14.6 | 34.4 | 65.1 | 14.7 |
| Visual Length | ScanQA (val) | SQA3D (test) | MSQA | |||
| EM@1 | EM@1-R | EM@1 | EM@1-R | EM@1 | EM@1-R | |
| tokens | 19.8 | 34.0 | 24.5 | 38.2 | 11.9 | 19.1 |
| tokens | 19.8 | 34.0 | 27.8 | 41.2 | 11.9 | 19.1 |
@@ 4.3 Performance on 3D Question Answering Tasks
Interestingly, we observe a slight performance drop when comparing SplatTalk with SplatTalk-ScanQA-FT. However, this drop becomes much more pronounced when compared with SlatTalk-3DVQA-FT. A plausible explanation is that fine-tuning on other QA datasets introduces overfitting, causing the model to adapt too specifically to those datasets. Furthermore, the observed performance drop suggests that MSR3D differs significantly from ScanQA and SQA3D, making direct knowledge transfer less effective. This distinction may also explain why LEO struggles to generalize in a zero-shot setting. Despite being trained on multiple QA datasets, none of them closely resemble MSR3D, limiting its ability to adapt without fine-tuning.
@@ 4.4 Ablation Studies
Gaussian Sampling Strategies. We explore various Gaussian sampling strategies for selecting tokens as inputs to the LLM, including: (1) Random Sampling—Gaussians are selected random across the scene. (2) Point Density Adaptive Sampling—Gaussians are sampled proportionally to the density of 3D points. (3) Furthest Point Sampling (FPS)—Gaussians are iteratively selected at locations maximally distant from previously chosen points. (4) Entropy Adaptive Sampling—Gaussians are prioritized based on language feature entropy, selecting points with the highest uncertainty.
We conduct experiments on a subset of the ScanQA validation set, consisting of 10 scenes. The results, presented in Table 4, demonstrate that entropy-adaptive sampling consistently outperforms all other sampling strategies. This finding highlights entropy-adaptive sampling as the most effective approach for selecting informative tokens, making it the optimal strategy for downstream 3D VQA tasks.
@@
A key observation from our results is that, apart from SQA3D, the performance on ScanQA and MSR3D remains unchanged regardless of the number of tokens used. Specifically, using only image-worth of high-entropy tokens achieves the same EM@1 score as using images-worth of tokens. This suggests that additional scene context does not contribute to performance improvements, which could be due to two reasons: 1) Dataset limitations—These datasets may inherently lack the need for additional scene or spatial information to improve performance; 2) Model limitations—The model may struggle to effectively utilize additional visual tokens for enhanced spatial understanding. In contrast, SQA3D shows a clear performance gain when incorporating more tokens, suggesting that this dataset genuinely requires richer 3D scene context to improve model understanding.
@@ 5 Conclusion
Appendix A Data Details
We train SplatTalk on 500 ScanNet scenes from the ScanQA training set. This stage does not involve any language supervisory signal and is purely self-supervised. We evaluate our method and baselines on 3 datasets: ScanQA [2] validation set, SQA3D [34] test set, and MSR3D [29] test set. The ScanQA validation set comprises 71 ScanNet scenes, while the SQA3D and MSR3D test sets includes 67 scenes, with significant overlap with ScanQA validation set.
Appendix B Training & Model Details
@@
where and are the ground-truth RGB image and the pseudo ground-truth language feature map respectively. and correspond to the predicted RGB image and language feature map respectively. We train SplatTalk on a one NVIDIA H100 (80GB) GPU.
@@ Appendix C Evaluation Metrics
On the other hand, EM@1 or Exact Match at 1, is a strict evaluation metric that only assigns a score if the generated response perfectly matches the reference answer, without any deviations. Even if the response is semantically correct but differs in phrasing, formatting, or minor details, the model is penalized. In contrast, EM@1-Refined offers a more flexible evaluation, allowing for slight variations in wording.
@@
Appendix D More Visualization on ScanQA

### v2

@@ Abstract
Language-guided 3D scene understanding is important for advancing applications in robotics, AR/VR, and human-computer interaction, enabling models to comprehend and interact with 3D environments through natural language. While 2D vision-language models (VLMs) have achieved remarkable success in 2D VQA tasks, progress in the 3D domain has been significantly slower due to the complexity of 3D data and the high cost of manual annotations. In this work, we introduce SplatTalk, a novel method that uses a generalizable 3D Gaussian Splatting (3DGS) framework to produce 3D tokens suitable for direct input into a pretrained LLM, enabling effective zero-shot 3D visual question answering (3D VQA) for scenes with only posed images. During experiments on multiple benchmarks, our approach outperforms both 3D models trained specifically for the task and previous 2D-LMM-based models utilizing only images (our setting), while achieving competitive performance with state-of-the-art 3D LMMs that additionally utilize 3D inputs. Project website: https://splat-talk.github.io/
@@ 1 Introduction
Previous works have explored integrating semantic features into NeRF [21] or 3D Gaussian Splatting (3DGS) [38], which rely only upon RGB image inputs. However, research on effectively associating 3D geometry with language fetures remains limited. Some methods align individual 3D locations with language features via implicit functions optimized individually for every scene [21, 10]. Others employ SAM [22] to first segment objects and object parts, followed by CLIP [39] to extract object-centric features from the generated masks. These approaches allow models to query information about individual points and/or objects. However, it disregards spatial relationships between objects, making it unsuitable for 3D VQA spatial reasoning tasks.
@@ 2 Related Work
Generalizable 3D Gaussian Splatting. This body of work explores generalizable 3DGS as an alternative to the per-scene optimization paradigm used in traditional 3DGS methods, aiming to improve novel view synthesis across unseen scenes. Our method builds on FreeSplat [46], which is designed for sparse-view synthesis in indoor scenes. The method begins by extracting multi-scale features from input RGB images using CNN backbones. It then constructs adaptive cost volumes from neighboring views to predict depth maps. These predicted depth maps are used to back-project 2D CNN-encoded features into 3D Gaussian triplets ( where denotes the Gaussian’s center, the weight, and the latent feature. To integrate multi-view information, FreeSplat introduces a Pixel-wise Triplet Fusion (PTF) module that uses a lightweight GRU to align and fuse overlapping 3D Gaussians across views efficiently. Finally, the fused triplets are passed through an MLP to decode rendering parameters, . Our approach extends the FreeSplat framework to enable the learning of generalizable 3D Gaussian language fields.
Similar to FreeSplat [46], [8, 5] leverage multi-view stereo (MVS) cues to refine depth estimation and improve geometric consistency in 3D Gaussian representations. These methods primarily focus on RGB novel view synthesis from sparse inputs ( views). Our work builds on the SOTA model, FreeSplat [46], but diverges significantly in both objectives and methodology. First, while previous methods focus on photorealistic RGB reconstruction, we aim at semantic 3D scene understanding. Second, instead of reconstructing only a localized portion of the scene as shown in these works, our goal is to represent entire 3D environments, capturing both spatial and semantic relationships holistically.
@@
ChatSplat [6] is a concurrent work that explores natural language conversation within the 3DGS framework. However, it optimizes on a per-scene basis and requires rendering 3D features into 2D feature maps during inference. Additionally, its training pipeline is not fully end-to-end, as it first trains the RGB Gaussians separately before incorporating language features. In contrast, our approach leverages scene latent features that encapsulate both RGB and semantic information, enabling the joint optimization of both modalities in a fully end-to-end manner. Our approach extends the FreeSplat framework to enable the learning of generalizable 3D Gaussian language fields.
@@ 3 Method
We provide an overview of our approach in Fig. 2, which consists of three main steps: 1) Training the feature autoencoder, 2) Self-supervised 3D-Language Gaussian Splatting Training and 3) 3D VQA Inference. First, we extract the high-dimensional language features from LLaVA-OV [23] for each RGB image by using the pretrained image encoder and multimodal projector. Since these features are high-dimensional, unbounded and sparse, we train an autoencoder to project them into a more compact, low-dimensional space. These low-dimensional features then serve as pseudo ground-truth for training the language representations in SplatTalk. SplatTalk architecture is a feed-forward network consisting of a Gaussian encoder and decoder that predicts Gaussian parameters for rendering. During 3D VQA inference stage, we extract the trained language features from the 3D Gaussians and directly feed them into the LLM, along with the language prompt, to generate free-form responses.
@@ 3.1 Integrating Language Features
[…]
@@
[…]
3.2 Extracting 3D Language Features
During inference, we directly extract language features at the mean position of each 3D Gaussian. Although the rendering process incorporates each Gaussian’s covariance and opacity when projecting 3D features to 2D, we demonstrate that scene semantics can be effectively captured by the mean feature of each Gaussian, , alone, without relying on these additional factors. This is supported by the conceptual similarity between our training optimization process and the Expectation-Maximization (EM) algorithm. For simplification, let be the loss, where is the predicted rendered feature map of frame at pixel and is the pseudo ground-truth feature map obtained from LLaVA-OV. Replacing by the rendering equation and solving for the optimal , we obtain where is the Gaussian’s contribution to pixel at viewpoint . This mirrors the EM procedure: estimating by rendering (E-step) and updating by minimizing reconstruction loss (M-step). Consequently, the optimal 3D feature for each Gaussian is a weighted sum of the 2D feature maps, and the set of 3D Gaussian features encodes the scene’s semantic holistically from 2D inputs. These features are then projected back into the original language embedding space using the pretrained decoder and directly provided as inputs to the LLM for downstream 3D-language tasks.
In open-vocabulary segmentation, querying Gaussian centers is effective [4, 38, 48] because each Gaussian is assumed to represent a distinct, meaningful object or region. These methods use object-centric embeddings (e.g., SAM-masked CLIP or LSeg) and retrieve objects via cosine similarity to text queries. In contrast, 3D reasoning tasks lack this one-to-one correspondence. Each Gaussian encodes parts of a scene-level representation, where meaning stems from spatial and relational structure. Querying single Gaussians is thus insufficient for context-dependent reasoning. While our method also queries at Gaussian means, this is not trivial; our EM-based formulation offers a principled justification.
3.3 3D VQA Inference
Entropy Adaptive Gaussian Sampling. SplatTalk leverages triplet fusion adopted in FreeSplat [46], which reduces the number of Gaussians. However, complex scenes still require significantly more Gaussians than the LLM’s visual token limit, and not all are equally informative. To address this, we propose entropy-based sampling: selecting the top Gaussians with the highest language feature entropy, where matches the LLM’s visual token capacity.
3D Tokens as Visual Inputs. After 3DGS training, each 3D Gaussian encodes a feature that is a weighted aggregation of 2D semantic information across all input views (Sec. 3.2). These features are also spatially grounded in their 3D coordinates. Because each token corresponds to a specific location in 3D space, its position is already implicitly encoded in the feature through the rendering and fusion process. Unlike 2D patches in a fixed grid, these 3D tokens form an unordered set, similar to a point cloud where spatial relationships are captured by the features themselves, not by sequence position. Therefore, we do not require explicit positional encodings or a fixed input order. The 3D tokens, extracted by sampling features at each Gaussian’s mean, are directly fed into the LLM.
@@ 4.3 Performance on 3D Question Answering Tasks
| Sampling Method | EM@1 | METEOR | ROUGE | SPICE | Avg |
| Random | 17.0 | 13.8 | 62.1 | 14.0 | 26.73 |
| Point Density | 16.2 | 14.0 | 59.7 | 14.1 | 26.00 |
| FPS | 17.7 | 13.8 | 62.0 | 13.6 | 26.78 |
| Entropy | 17.1 | 14.6 | 61.7 | 14.2 | 26.90 |
| Visual Length | ScanQA (val) | SQA3D (test) | MSR3D (test) | |||
| EM@1 | EM@1-R | EM@1 | EM@1-R | EM@1 | EM@1-R | |
| tokens | 16.2 | 30.7 | 25.8 | 38.0 | 7.9 | 17.7 |
| tokens | 17.1 | 32.2 | 26.1 | 41.3 | 14.1 | 23.3 |
@@
Interestingly, we observe a slight performance drop when comparing SplatTalk with SplatTalk-ScanQA-FT. However, this drop becomes much more pronounced when compared with SplatTalk-3DVQA-FT. A plausible explanation is that fine-tuning on other QA datasets introduces overfitting, causing the model to adapt too specifically to those datasets. Furthermore, the observed performance drop suggests that MSR3D differs significantly from ScanQA and SQA3D, making direct knowledge transfer less effective. This distinction may also explain why LEO struggles to generalize in a zero-shot setting. Despite being trained on multiple QA datasets, none of them closely resemble MSR3D, limiting its ability to adapt without fine-tuning.
@@ 4.4 Ablation Studies
Gaussian Sampling Strategies. We explore various Gaussian sampling strategies for selecting tokens as inputs to the LLM, including: (1) Random Sampling—Gaussians are selected randomly across the scene. (2) Point Density Adaptive Sampling—Gaussians are sampled proportionally to the density of 3D points. (3) Furthest Point Sampling (FPS)—Gaussians are iteratively selected at locations maximally distant from previously chosen points. (4) Entropy Adaptive Sampling—Gaussians are prioritized based on language feature entropy, selecting points with the highest uncertainty.
We conduct experiments on the ScanQA validation set. The results, presented in Table 4, demonstrate that entropy-adaptive sampling outperforms other sampling strategies, followed by FPS. This finding highlights entropy-adaptive sampling as the most effective approach for selecting informative tokens, making it the optimal strategy for downstream 3D VQA tasks.
@@
A key observation is that performance improved across all datasets as the number of input tokens increased, with MSR3D showing the largest gain, nearly doubling in EM@1. In contrast, ScanQA and SQA3D saw only marginal improvements, even with significantly more tokens. This suggests that additional scene context may not substantially improve performance, possibly due to: (1) Dataset limitations where the tasks may not require rich spatial context; or (2) Model limitations where the model may be unable to effectively leverage extra visual tokens for spatial reasoning.
@@ 5 Conclusion
Acknowledgment
We thank the creators of the datasets and codebases used in our work, ScanQA, SQA3D, MSR3D, ScanNet, LLaVA-OV, and FreeSplat, for making their resources publicly available. We are also grateful to the anonymous reviewers for their constructive feedback during the review process. Special thanks to Xiang Li and James M. Rehg for the insightful discussions.
Supplementary Material
6 Data Details
We train SplatTalk on 500 ScanNet scenes from the ScanQA training set. This stage does not involve any language supervisory signal and is purely self-supervised. We evaluate our method and baselines on 3 datasets: ScanQA [2] validation set, SQA3D [34] test set, and MSR3D [29] test set. The ScanQA validation set comprises 71 ScanNet scenes, while the SQA3D and MSR3D test sets include 67 scenes, with significant overlap with ScanQA validation set.
7 Training & Model Details
@@
where and are the ground-truth RGB image and the pseudo ground-truth language feature map respectively. and correspond to the predicted RGB image and language feature map respectively. We train SplatTalk on one NVIDIA H100 (80GB) GPU.
@@ 8 Evaluation Metrics
On the other hand, EM@1 or Exact Match at top-1, is a strict evaluation metric that only assigns a score if the generated response perfectly matches the reference answer, without any deviations. Even if the response is semantically correct but differs in phrasing, formatting, or minor details, the model is penalized. In contrast, EM@1-Refined offers a more flexible evaluation, allowing for slight variations in wording.
@@
9 More Visualization on ScanQA